# Zooming In on Fakes: A Novel Dataset for Localized AI-Generated Image Detection with Forgery Amplification Approach

### v3

@@ Introduction
The rapid advancement of generative models, such as Generative Adversarial Networks (GANs) (DBLP:conf/iclr/KarrasALL18) and Diffusion Models (DMs) (dangdiff), enable fine-grained image modifications through simple user interactions like masks, and prompts (ju2024brushnet). These techniques raise serious concerns about image authenticity (ferreira2020review), especially on social media platforms, while democratizing the creation of creative content. Consequently, the ability to detect whether visual content has been altered is becoming increasingly critical.
Around the AI-generated content (AIGC), a few early researches construct various datasets (wang2023dire; zhu2023genimage) and benchmarks (wang2020cnn; ojha2023towards) to improve the detection performance on fully synthesized images. However, these efforts often overlook localized generation scenarios, where only specific regions are modified. Although some recent works (guillaro2023trufor; he2025survey) have attempted to localize manipulations, their progress is constrained by limitations in existing datasets and detection paradigm.
Figure 1: Comparison of four forgery scenarios in existing datasets. They mainly cover full-generated images and object-level forgeries, while forgeries in stuff and background regions remain largely unaddressed. Red regions show ground-truth forgeries. State-of-the-art models (FatFormer (liu2024forgery) and SparseViT (su2024can)) struggle with these new cases. Our proposed NFA-ViT achieves robust detection across all four scenarios. The source data comes from open-source datasets GRE (sun2024rethinking), COCO (lin2014microsoft), and ImageNet (deng2009imagenet).
@@
NIST16  (guan2019mfc) 0 564 1 - Small
DEFACTO  (mahfoudi2019defacto) - 149,000 1 - Small
IMD2020  (novozamsky2020imd2020) 35,000 35,000 1 - Small
CocoGLIDE  (guillaro2023trufor) 512 512 - 1 Small
AutoSplice  (jia2023autosplice) 2,273 3,621 - 1 Small/Medium
TGIF  (mareen2024tgif) 3,124 74,976 - 2 Small
GRE  (sun2024rethinking) - 228,650 2 3 Small
SID-Set  (huang2025sidasocialmediaimage) 70,000 140,000 - 2 Small/Full
@@
Specifically, current localized AIGC datasets (guillaro2023trufor) suffer from two major limitations. (1) Pervasive forgery region bias. Previous datasets focus on salient objects or synthetic rectangular patches while defining forged areas, neglecting complex scene-level elements like sky, ground, vegetation, or structural background. Detectors trained on such data tend to overfit to object-centric artifacts and fail to generalize to more subtle or spatially distributed forgeries. (2) Uncontrollable editing quality, which further limits their effectiveness. Many generated samples exhibit unrealistic textures, compression artifacts, or visible boundary seams due to low-quality generation pipelines and a lack of quality control. These flaws not only reduce visual plausibility but also make detection easier, masking the true difficulty of localized forgery detection in real-world scenarios. Fig. 1 illustrates how data limitations affect model detection. The model fails to detect when faced with out-of-distribution data or complex scene-level elements.
To address these challenges, we present the Broader Region Generation (BR-Gen) dataset, a large-scale and high-quality benchmark containing 150,000 locally forged images with diverse region coverage. BR-Gen targets underrepresented “stuff” and “background” categories—including sky, ground, wall, grass, and vegetation—substantially broadening the scope of localized forgeries beyond objects. The dataset is constructed through a fully automated “Perception-Creation-Evaluation” pipeline that ensures semantic integrity and visual realism. Specifically, we use grounding and segmentation models to guide localized editing (ravi2024sam2; liu2023grounding), diffusion-based generative models for content synthesis (podell2023sdxl; zhuang2024task; wang2025training), and multi-stage perceptual evaluation metrics (fu2023dreamsim; radford2021learning) to validate image quality. Compared to prior datasets, BR-Gen offers broader region diversity, more realistic forgeries, and stronger alignment with real-world editing patterns.
Despite recent advances (su2024can; guillaro2023trufor; liu2022pscc) in localized forgery detection, existing methods exhibit two major limitations when evaluated on BR-Gen. First, they tend to detect generic discrepancies between regions without the capability to identify which parts are truly forged. This often leads to overfitting to dataset-specific biases, such as the typical size, shape, or location of forged areas, and results in significant errors when these patterns vary. See Case 3 in Fig.5 for an example. Second, when forgeries are small or embedded in visually inconspicuous regions (e.g., within complex backgrounds or “stuff” areas), the manipulation signals become extremely weak and are easily overshadowed by surrounding authentic content, rendering them difficult to detect. This issue is illustrated in Case 4 of Fig.5.
@@
Figure 2: The automated pipeline for the BR-Gen dataset consists of three iterative stages: Perception, Creation, and Evaluation. These stages are applied to produce high-quality localized generation datasets through progressive refinement. All samples are sourced from publicly available datasets (zhou2017places; lin2014microsoft; deng2009imagenet).
@@
In recent years, Artificial Intelligence Generated Content (AIGC) has become widely used. Datasets (wang2020cnn; ojha2023towards) containing both real and generated images have been organized for training and evaluating detection systems. CNNSpot (wang2020cnn) uses GAN-generated images (DBLP:conf/iclr/KarrasALL18; chen2024decoupled), and Chameleon (yan2024sanity) provides highly realistic test cases. However, these support only image-level tasks. For localized detection, existing datasets (Tab. 1) often use object masks from COCO (lin2014microsoft) or SAM (kirillov2023segment), focusing on countable objects. This ignores large regions like sky or background, leading to bias and poor model generalization on such areas.
@@
The need for detecting AI-generated content has existed since the rise of deep learning. Early detection used spatial cues like color and reflection (mccloskey2018detecting; o2012exposing). As GANs improved, methods like CNNSpot (wang2020cnn) used data augmentation for better generalization (karras2017progressive; chen2025dual; yang2025all; lin2025seeing). Recent works exploit frequency or reconstruction patterns (wang2023dire; qian2020thinking), but still focus on whole images. For localization, ManTra-Net (wu2019mantra) uses LSTM analysis, Trufor (guillaro2023trufor) uses noise patterns, and SparseViT (su2024can) uses sparse attention for top performance. Yet, they fail on complex or non-object regions, highlighting the need for our approach.
@@
Recent datasets for detecting localized forgery based on generative models (sun2024rethinking; guillaro2023trufor) have emerged. To address the gaps in existing datasets, we have taken into account the the neglected local edits in Stuff and Background, proposing a high-quality, scene-based local generation dataset named the Broader Region Generation (BR-Gen). We propose an automated pipeline with open-source models (ravi2024sam2; liu2023grounding; DBLP:journals/corr/abs-2502-13923), generating local edited images from unannotated ones. As shown in Tab. 1, BR-Gen takes into full account diverse generation methods and tampering areas, addressing the shortcomings in the types of previous datasets.
@@
We sampled images from three large-scale datasets like previous works (jia2023autosplice): ImageNet (deng2009imagenet), COCO (lin2014microsoft), and Places (zhou2017places). These datasets provide diverse scenes with rich semantic content, enhancing the diversity of dimensions.
@@
To simulate real-world image editing processes while maintaining content and semantic consistency, we designed an automated pipeline that integrates multiple open-source models (ravi2024sam2; liu2023grounding; DBLP:journals/corr/abs-2502-13923), as illustrated in Fig. 2. The pipeline comprises three stages: “Perception-Creation-Evaluation”. Below, we briefly describe each stage, with detailed process descriptions and dataset statistics provided in the appendix.
@@
This stage identifies regions of interest in real images and extracts semantic information to guide subsequent creation stages. We select the candidate categories by “thing category” and “stuff category” from COCO. GroundingDINO (liu2023grounding) detects object bounding boxes for these categories, which are then converted into masks by SAM2 (ravi2024sam2). “Stuff” masks are obtained directly, while “background” masks are derived by inverting “thing” masks. To balance category distribution, we control the number of instances per category.
For prompt-based inpainting, we use Qwen2.5-VL (DBLP:journals/corr/abs-2502-13923) to generate image descriptions. Then, we apply Semantic Perturbation to increase semantic diversity: for “stuff”, we replace content within “#” symbols (e.g., “the blue sky” $\to$ “the starry sky”); for “background”, we replace text outside “\$” symbols. Extra semantically similar object lists ensure plausible edits.
Figure 3: Partial examples from BR-Gen dataset. The real data comes from the open datasets Places (zhou2017places), COCO (lin2014microsoft), and ImageNet (deng2009imagenet).
@@
This stage uses region masks and guiding prompts to generate locally forged images. We employ five diversity state-of-the-art inpainting methods: two GAN-based (LaMa (suvorov2022resolution), MAT (li2022mat)) and three diffusion-based (SDXL (podell2023sdxl), BrushNet (ju2024brushnet), PowerPaint (zhuang2024task)). GAN methods require only the image and mask, while diffusion methods also require the guiding prompt.
@@
This stage filters high-quality generated images using image quality assessment methods. Image quality is assessed based on: (a) BRISQUE (mittal2012no) for structural integrity; (b) DreamSim (fu2023dreamsim) for image similarity; and (c) CLIP scores for prompt alignment. Low-quality samples are removed. Missing data is replenished through iterative generation to maintain dataset size.
@@
To further enhance the performance of local AIGC detection, we propose the Noise-guided Forgery Amplification Vision Transformer (NFA-ViT) to take advantages of the non-homologous between the generated regions and real regions. Noise fingerprints and related features exhibit discrepancies across different regions (guillaro2023trufor; wang2023dire). However, related studies suffer from significant limitations, such as cannot determine which region is forged and being insensitive to subtle and tiny-area forgeries. We propose a novel detection mechanism called forgery amplification. NFA-ViT leverages noise information as guidance to amplify and diffuse localized forgery features across the entire image, making forgery features more distinguishable while ensuring that the judgment of real images remains unaffected.
Fig. 4(a) illustrates the overall framework. For an input RGB image $x$, we first use the noise extractor Noiseprint++ (guillaro2023trufor) to extract the noise trace $n$ of the image. We verify the effectiveness of the features and provide detailed information in the appendix. Subsequently, $x$ and $n$ are jointly fed into a dual-branch network. For each stage, the Noise-guided Mask $M_{i}^{n}$ from the noise branch is used to guide the learning of the proposed Noise-guided Amplification Attention (NAA) in the image branch. With the NAA, real regions directly focus on the differential forgery regions, gradually diffusing forgery features into real regions:
@@
For simplicity, we describe the attention workflow for a single attention head. We first introduce Fix-Sparse Attention (su2024can), which helps refine localization by eliminating irrelevant semantic information. However, Fix-Sparse Attention disrupts semantics from a global perspective to learn non-semantic features, lacking the ability to aggregate and recognize local information.
@@
Localized Detection ManTranet (wu2019mantra) 0.822 0.123 - 0.069 0.008 ($\downarrow$ 0.133) 0.074 0.061 0.077 0.058
MVSS-Net (dong2022mvss) 0.862 0.183 0.344 0.122 0.029 ($\downarrow$ 0.424) 0.154 0.098 0.154 0.092
PSCC-Net (liu2022pscc) 0.806 0.253 0.284 0.164 0.052 ($\downarrow$ 0.426) 0.166 0.161 0.170 0.155
Trufor (guillaro2023trufor) 0.881 0.295 0.319 0.194 0.048 ($\downarrow$ 0.630) 0.195 0.194 0.199 0.187
SparseViT (su2024can) 0.735 0.277 - 0.203 0.049 ($\downarrow$ 0.649) 0.214 0.186 0.205 0.202
AIGC Detection LGrad (tan2023learning) 0.974 0.165 0.635 0.088 - 0.101 ($\downarrow$ 0.762) 0.057 0.093 0.085
FreqNet (DBLP:conf/aaai/Tan0WGLW24) 0.767 0.360 0.472 0.231 - 0.244 ($\downarrow$ 0.671) 0.228 0.240 0.229
NPR (tan2024rethinking) 0.894 0.443 0.501 0.300 - 0.323 ($\downarrow$ 0.602) 0.290 ($\downarrow$ 0.662) 0.318 0.289
FatFormer (liu2024forgery) 0.989 0.493 0.606 0.331 - 0.358 ($\downarrow$ 0.626) 0.321 ($\downarrow$ 0.629) 0.349 0.310
Table 2: The cross-domain results on BR-Gen. The evaluation methods include AIGC detection and localization detection. We bold the best result and mark the second-best result with an underline. The red decline values indicate the level of performance decrease compared to the original dataset (ma2025imdl; liu2024forgery; tan2024rethinking). Since some methods don’t provide corresponding indicators, some values are missing.
@@
Localized Detection ManTranet (wu2019mantra) 0.853 0.774 - 0.760 0.632 0.772 0.760 0.778 0.754
MVSS-Net (dong2022mvss) 0.903 0.892 0.924 0.883 0.671 0.913 0.846 0.889 0.856
PSCC-Net (liu2022pscc) 0.935 0.894 0.937 0.861 0.705 0.898 0.840 0.867 0.844
Trufor (guillaro2023trufor) 0.944 0.918 0.942 0.896 0.779 0.915 0.865 0.903 0.871
SparseViT (su2024can) 0.984 0.946 - 0.911 0.824 0.958 0.872 0.931 0.907
AIGC Detection LGrad (tan2023learning) 0.937 0.831 0.872 0.755 - 0.801 0.732 0.775 0.738
DIRE (wang2023dire) 0.939 0.823 0.825 0.742 - 0.750 0.744 0.762 0.739
FreqNet (DBLP:conf/aaai/Tan0WGLW24) 0.825 0.699 0.702 0.631 - 0.659 0.614 0.648 0.622
NPR (tan2024rethinking) 0.946 0.922 0.933 0.902 - 0.938 0.884 0.921 0.893
FatFormer (liu2024forgery) 0.990 0.961 0.971 0.935 - 0.955 0.913 0.949 0.915
AIDE (yan2024sanity) 0.986 0.964 0.973 0.941 - 0.950 0.932 0.965 0.937
@@
Current multi-level feature fusion methods often use addition or concatenation (lin2017feature), producing feature maps through fixed linear aggregation without accounting for the varying contributions of hierarchical features to final maps. To improve region mask prediction, we propose a simple yet efficient decoder design. This approach introduces learnable scaling parameters $\gamma_{i}\ (1\leq i\leq 4)$ for each hierarchical feature map, enabling adaptive weighted fusion by modulating layer-wise contributions.
@@
For detection tasks, we use a lightweight backbone network (he2016deep) to extract features $F_{4}$, which generate predictions $\hat{y}$. For localization tasks, predicted masks $\hat{M}$ from the Weighted Decoder are utilized. Given ground truth labels $y$ and ground truth region masks $M$, the NFA-ViT model is trained using the following objective function:
@@
Implementation Details. For NFA-ViT, we use SegFormer (xie2021segformer) as the backbone, with the image and noise encoders being the b2 and b0 versions, respectively. During training, the model is optimized using the Adam optimizer (kinga2015method), with an initial learning rate of $5\times 10^{-3}$ and a weight decay of $1\times 10^{-6}$. Using the Warmup and CosineAnnealing to help models achieve better convergence. In the Noise-Guided Amplification Attention mechanism, the $\text{Top-}k$ ratio is set to 25%. Equal weights are given to all parts of the loss function. All experiments are run for 30 epochs with a batch size of 64.
@@
Cross-domain on BR-Gen. To evaluate data bias in current generated image detection tasks, we conducted cross-domain testing on the BR-Gen dataset using detection models from two tasks: AIGC detection and local AIGC detection. We directly test the released trained models of those models, which are trained on data (novozamsky2020imd2020; karras2017progressive) that shares the same source as BR-Gen. We divided the BR-Gen’s into “Split A” and “Split B” according generation method and mask sources. Among them, “Split A” is categorized based on the generation method, specifically GAN and diffusion, while “Split B” is categorized based on the source of the masks, i.e., stuff and background. For each model, we also compared its performance drop relative to its original report (ma2025imdl; liu2024forgery; tan2024rethinking), under the corresponding data distribution. The experimental results are shown in Tab. 2.
@@
Figure 5: Localization results of different models. We compared images generated by two types of masks. All samples are sourced from publicly available datasets (zhou2017places; lin2014microsoft; deng2009imagenet).

### v4

@@ Introduction
The rapid advancement of generative models, such as Generative Adversarial Networks (GANs) (Karras et al. 2018a) and Diffusion Models (DMs) (Dang et al. 2025), enable fine-grained image modifications through simple user interactions like masks, and prompts (Ju et al. 2024). These techniques raise serious concerns about image authenticity (Ferreira et al. 2020), especially on social media platforms, while democratizing the creation of creative content. Consequently, the ability to detect whether visual content has been altered is becoming increasingly critical.
Around the AI-generated content (AIGC), a few early researches construct various datasets (Wang et al. 2023; Zhu et al. 2023) and benchmarks (Wang et al. 2020; Ojha et al. 2023) to improve the detection performance on fully synthesized images. However, these efforts often overlook localized generation scenarios, where only specific regions are modified. Although some recent works (Guillaro et al. 2023; He et al. 2025) have attempted to localize manipulations, their progress is constrained by limitations in existing datasets and detection paradigm.
Figure 1: Comparison of four forgery scenarios in existing datasets. They mainly cover full-generated images and object-level forgeries, while forgeries in stuff and background regions remain largely unaddressed. Red regions show ground-truth forgeries. State-of-the-art models (FatFormer (Liu et al. 2024) and SparseViT (Su et al. 2025)) struggle with these new cases. Our proposed NFA-ViT achieves robust detection across all four scenarios. The source data comes from open-source datasets GRE (Sun et al. 2024), COCO (Lin et al. 2014), and ImageNet (Deng et al. 2009).
@@
NIST16  (Guan et al. 2019) 0 564 1 - Small
DEFACTO  (Mahfoudi et al. 2019) - 149,000 1 - Small
IMD2020  (Novozamsky et al. 2020) 35,000 35,000 1 - Small
CocoGLIDE  (Guillaro et al. 2023) 512 512 - 1 Small
AutoSplice  (Jia et al. 2023) 2,273 3,621 - 1 Small/Medium
TGIF  (Mareen et al. 2024) 3,124 74,976 - 2 Small
GRE  (Sun et al. 2024) - 228,650 2 3 Small
SID-Set  (Huang et al. 2025) 70,000 140,000 - 2 Small/Full
@@
Specifically, current localized AIGC datasets (Guillaro et al. 2023) suffer from two major limitations. (1) Pervasive forgery region bias. Previous datasets focus on salient objects or synthetic rectangular patches while defining forged areas, neglecting complex scene-level elements like sky, ground, vegetation, or structural background. Detectors trained on such data tend to overfit to object-centric artifacts and fail to generalize to more subtle or spatially distributed forgeries. (2) Uncontrollable editing quality, which further limits their effectiveness. Many generated samples exhibit unrealistic textures, compression artifacts, or visible boundary seams due to low-quality generation pipelines and a lack of quality control. These flaws not only reduce visual plausibility but also make detection easier, masking the true difficulty of localized forgery detection in real-world scenarios. Fig. 1 illustrates how data limitations affect model detection. The model fails to detect when faced with out-of-distribution data or complex scene-level elements.
To address these challenges, we present the Broader Region Generation (BR-Gen) dataset, a large-scale and high-quality benchmark containing 150,000 locally forged images with diverse region coverage. BR-Gen targets underrepresented “stuff” and “background” categories—including sky, ground, wall, grass, and vegetation—substantially broadening the scope of localized forgeries beyond objects. The dataset is constructed through a fully automated “Perception-Creation-Evaluation” pipeline that ensures semantic integrity and visual realism. Specifically, we use grounding and segmentation models to guide localized editing (Ravi et al. 2024; Liu et al. 2023), diffusion-based generative models for content synthesis (Podell et al. 2023; Zhuang et al. 2024; Wang et al. 2025), and multi-stage perceptual evaluation metrics (Fu et al. 2023; Radford et al. 2021) to validate image quality. Compared to prior datasets, BR-Gen offers broader region diversity, more realistic forgeries, and stronger alignment with real-world editing patterns.
Despite recent advances (Su et al. 2025; Guillaro et al. 2023; Liu et al. 2022b) in localized forgery detection, existing methods exhibit two major limitations when evaluated on BR-Gen. First, they tend to detect generic discrepancies between regions without the capability to identify which parts are truly forged. This often leads to overfitting to dataset-specific biases, such as the typical size, shape, or location of forged areas, and results in significant errors when these patterns vary. See Case 3 in Fig.5 for an example. Second, when forgeries are small or embedded in visually inconspicuous regions (e.g., within complex backgrounds or “stuff” areas), the manipulation signals become extremely weak and are easily overshadowed by surrounding authentic content, rendering them difficult to detect. This issue is illustrated in Case 4 of Fig.5.
@@
Figure 2: The automated pipeline for the BR-Gen dataset consists of three iterative stages: Perception, Creation, and Evaluation. These stages are applied to produce high-quality localized generation datasets through progressive refinement. All samples are sourced from publicly available datasets (Zhou et al. 2017; Lin et al. 2014; Deng et al. 2009).
@@
In recent years, Artificial Intelligence Generated Content (AIGC) has become widely used. Datasets (Wang et al. 2020; Ojha et al. 2023) containing both real and generated images have been organized for training and evaluating detection systems. CNNSpot (Wang et al. 2020) uses GAN-generated images (Karras et al. 2018a; Chen et al. 2024), and Chameleon (Yan et al. 2025) provides highly realistic test cases. However, these support only image-level tasks. For localized detection, existing datasets (Tab. 6) often use object masks from COCO (Lin et al. 2014) or SAM (Kirillov et al. 2023), focusing on countable objects. This ignores large regions like sky or background, leading to bias and poor model generalization on such areas.
@@
The need for detecting AI-generated content has existed since the rise of deep learning. Early detection used spatial cues like color and reflection (McCloskey and Albright 2018; O’brien and Farid 2012). As GANs improved, methods like CNNSpot (Wang et al. 2020) used data augmentation for better generalization (Karras et al. 2018b; Chen et al. 2025; Yang et al. 2025; Lin et al. 2025). Recent works exploit frequency or reconstruction patterns (Wang et al. 2023; Qian et al. 2020), but still focus on whole images. For localization, ManTra-Net (Wu et al. 2019) uses LSTM analysis, Trufor (Guillaro et al. 2023) uses noise patterns, and SparseViT (Su et al. 2025) uses sparse attention for top performance. Yet, they fail on complex or non-object regions, highlighting the need for our approach.
@@
Recent datasets for detecting localized forgery based on generative models (Sun et al. 2024; Guillaro et al. 2023) have emerged. To address the gaps in existing datasets, we have taken into account the the neglected local edits in Stuff and Background, proposing a high-quality, scene-based local generation dataset named the Broader Region Generation (BR-Gen). We propose an automated pipeline with open-source models (Ravi et al. 2024; Liu et al. 2023; Bai et al. 2025), generating local edited images from unannotated ones. As shown in Tab. 6, BR-Gen takes into full account diverse generation methods and tampering areas, addressing the shortcomings in the types of previous datasets.
@@
We sampled images from three large-scale datasets like previous works (Jia et al. 2023): ImageNet (Deng et al. 2009), COCO (Lin et al. 2014), and Places (Zhou et al. 2017). These datasets provide diverse scenes with rich semantic content, enhancing the diversity of dimensions.
@@
To simulate real-world image editing processes while maintaining content and semantic consistency, we designed an automated pipeline that integrates multiple open-source models (Ravi et al. 2024; Liu et al. 2023; Bai et al. 2025), as illustrated in Fig. 7. The pipeline comprises three stages: “Perception-Creation-Evaluation”. Below, we briefly describe each stage, with detailed process descriptions and dataset statistics provided in the appendix.
@@
This stage identifies regions of interest in real images and extracts semantic information to guide subsequent creation stages. We select the candidate categories by “thing category” and “stuff category” from COCO. GroundingDINO (Liu et al. 2023) detects object bounding boxes for these categories, which are then converted into masks by SAM2 (Ravi et al. 2024). “Stuff” masks are obtained directly, while “background” masks are derived by inverting “thing” masks. To balance category distribution, we control the number of instances per category.
For prompt-based inpainting, we use Qwen2.5-VL (Bai et al. 2025) to generate image descriptions. Then, we apply Semantic Perturbation to increase semantic diversity: for “stuff”, we replace content within “#” symbols (e.g., “the blue sky” $\to$ “the starry sky”); for “background”, we replace text outside “\$” symbols. Extra semantically similar object lists ensure plausible edits.
Figure 3: Partial examples from BR-Gen dataset. The real data comes from the open datasets Places (Zhou et al. 2017), COCO (Lin et al. 2014), and ImageNet (Deng et al. 2009).
@@
This stage uses region masks and guiding prompts to generate locally forged images. We employ five diversity state-of-the-art inpainting methods: two GAN-based (LaMa (Suvorov et al. 2022), MAT (Li et al. 2022)) and three diffusion-based (SDXL (Podell et al. 2023), BrushNet (Ju et al. 2024), PowerPaint (Zhuang et al. 2024)). GAN methods require only the image and mask, while diffusion methods also require the guiding prompt.
@@
This stage filters high-quality generated images using image quality assessment methods. Image quality is assessed based on: (a) BRISQUE (Mittal et al. 2012) for structural integrity; (b) DreamSim (Fu et al. 2023) for image similarity; and (c) CLIP scores for prompt alignment. Low-quality samples are removed. Missing data is replenished through iterative generation to maintain dataset size.
@@
To further enhance the performance of local AIGC detection, we propose the Noise-guided Forgery Amplification Vision Transformer (NFA-ViT) to take advantages of the non-homologous between the generated regions and real regions. Noise fingerprints and related features exhibit discrepancies across different regions (Guillaro et al. 2023; Wang et al. 2023). However, related studies suffer from significant limitations, such as cannot determine which region is forged and being insensitive to subtle and tiny-area forgeries. We propose a novel detection mechanism called forgery amplification. NFA-ViT leverages noise information as guidance to amplify and diffuse localized forgery features across the entire image, making forgery features more distinguishable while ensuring that the judgment of real images remains unaffected.
Fig. 4(a) illustrates the overall framework. For an input RGB image $x$, we first use the noise extractor Noiseprint++ (Guillaro et al. 2023) to extract the noise trace $n$ of the image. We verify the effectiveness of the features and provide detailed information in the appendix. Subsequently, $x$ and $n$ are jointly fed into a dual-branch network. For each stage, the Noise-guided Mask $M_{i}^{n}$ from the noise branch is used to guide the learning of the proposed Noise-guided Amplification Attention (NAA) in the image branch. With the NAA, real regions directly focus on the differential forgery regions, gradually diffusing forgery features into real regions:
@@
For simplicity, we describe the attention workflow for a single attention head. We first introduce Fix-Sparse Attention (Su et al. 2025), which helps refine localization by eliminating irrelevant semantic information. However, Fix-Sparse Attention disrupts semantics from a global perspective to learn non-semantic features, lacking the ability to aggregate and recognize local information.
@@
Localized Detection ManTranet (Wu et al. 2019) 0.822 0.123 - 0.069 0.008 ($\downarrow$ 0.133) 0.074 0.061 0.077 0.058
MVSS-Net (Dong et al. 2022) 0.862 0.183 0.344 0.122 0.029 ($\downarrow$ 0.424) 0.154 0.098 0.154 0.092
PSCC-Net (Liu et al. 2022b) 0.806 0.253 0.284 0.164 0.052 ($\downarrow$ 0.426) 0.166 0.161 0.170 0.155
Trufor (Guillaro et al. 2023) 0.881 0.295 0.319 0.194 0.048 ($\downarrow$ 0.630) 0.195 0.194 0.199 0.187
SparseViT (Su et al. 2025) 0.735 0.277 - 0.203 0.049 ($\downarrow$ 0.649) 0.214 0.186 0.205 0.202
AIGC Detection LGrad (Tan et al. 2023) 0.974 0.165 0.635 0.088 - 0.101 ($\downarrow$ 0.762) 0.057 0.093 0.085
FreqNet (Tan et al. 2024a) 0.767 0.360 0.472 0.231 - 0.244 ($\downarrow$ 0.671) 0.228 0.240 0.229
NPR (Tan et al. 2024b) 0.894 0.443 0.501 0.300 - 0.323 ($\downarrow$ 0.602) 0.290 ($\downarrow$ 0.662) 0.318 0.289
FatFormer (Liu et al. 2024) 0.989 0.493 0.606 0.331 - 0.358 ($\downarrow$ 0.626) 0.321 ($\downarrow$ 0.629) 0.349 0.310
Table 2: The cross-domain results on BR-Gen. The evaluation methods include AIGC detection and localization detection. We bold the best result and mark the second-best result with an underline. The red decline values indicate the level of performance decrease compared to the original dataset (Ma et al. 2025; Liu et al. 2024; Tan et al. 2024b). Since some methods don’t provide corresponding indicators, some values are missing.
@@
Localized Detection ManTranet (Wu et al. 2019) 0.853 0.774 - 0.760 0.632 0.772 0.760 0.778 0.754
MVSS-Net (Dong et al. 2022) 0.903 0.892 0.924 0.883 0.671 0.913 0.846 0.889 0.856
PSCC-Net (Liu et al. 2022b) 0.935 0.894 0.937 0.861 0.705 0.898 0.840 0.867 0.844
Trufor (Guillaro et al. 2023) 0.944 0.918 0.942 0.896 0.779 0.915 0.865 0.903 0.871
SparseViT (Su et al. 2025) 0.984 0.946 - 0.911 0.824 0.958 0.872 0.931 0.907
AIGC Detection LGrad (Tan et al. 2023) 0.937 0.831 0.872 0.755 - 0.801 0.732 0.775 0.738
DIRE (Wang et al. 2023) 0.939 0.823 0.825 0.742 - 0.750 0.744 0.762 0.739
FreqNet (Tan et al. 2024a) 0.825 0.699 0.702 0.631 - 0.659 0.614 0.648 0.622
NPR (Tan et al. 2024b) 0.946 0.922 0.933 0.902 - 0.938 0.884 0.921 0.893
FatFormer (Liu et al. 2024) 0.990 0.961 0.971 0.935 - 0.955 0.913 0.949 0.915
AIDE (Yan et al. 2025) 0.986 0.964 0.973 0.941 - 0.950 0.932 0.965 0.937
@@
Current multi-level feature fusion methods often use addition or concatenation (Lin et al. 2017), producing feature maps through fixed linear aggregation without accounting for the varying contributions of hierarchical features to final maps. To improve region mask prediction, we propose a simple yet efficient decoder design. This approach introduces learnable scaling parameters $\gamma_{i}\ (1\leq i\leq 4)$ for each hierarchical feature map, enabling adaptive weighted fusion by modulating layer-wise contributions.
@@
For detection tasks, we use a lightweight backbone network (He et al. 2016) to extract features $F_{4}$, which generate predictions $\hat{y}$. For localization tasks, predicted masks $\hat{M}$ from the Weighted Decoder are utilized. Given ground truth labels $y$ and ground truth region masks $M$, the NFA-ViT model is trained using the following objective function:
@@
Implementation Details. For NFA-ViT, we use SegFormer (Xie et al. 2021) as the backbone, with the image and noise encoders being the b2 and b0 versions, respectively. During training, the model is optimized using the Adam optimizer (Kinga et al. 2015), with an initial learning rate of $5\times 10^{-3}$ and a weight decay of $1\times 10^{-6}$. Using the Warmup and CosineAnnealing to help models achieve better convergence. In the Noise-Guided Amplification Attention mechanism, the $\text{Top-}k$ ratio is set to 25%. Equal weights are given to all parts of the loss function. All experiments are run for 30 epochs with a batch size of 64.
@@
Cross-domain on BR-Gen. To evaluate data bias in current generated image detection tasks, we conducted cross-domain testing on the BR-Gen dataset using detection models from two tasks: AIGC detection and local AIGC detection. We directly test the released trained models of those models, which are trained on data (Novozamsky et al. 2020; Karras et al. 2018b) that shares the same source as BR-Gen. We divided the BR-Gen’s into “Split A” and “Split B” according generation method and mask sources. Among them, “Split A” is categorized based on the generation method, specifically GAN and diffusion, while “Split B” is categorized based on the source of the masks, i.e., stuff and background. For each model, we also compared its performance drop relative to its original report (Ma et al. 2025; Liu et al. 2024; Tan et al. 2024b), under the corresponding data distribution. The experimental results are shown in Tab. 2.
@@
Figure 5: Localization results of different models. We compared images generated by two types of masks. All samples are sourced from publicly available datasets (Zhou et al. 2017; Lin et al. 2014; Deng et al. 2009).